I have gotten the answer incorrect three times. Can you please assist in answering this question:

Two methods can be used to produce solar panels for electric power generation. Method 1 will have an initial cost of $520,000, an AOC of $230,000 per year, and $155,000 salvage value after its 3-year life. Method 2 will cost $950,000 with an AOC of $155,000 and a $210,000 salvage value after its 5-year life. Assume your boss asked you to determine which method is better, but she wants the analysis done over a three-year planning period. You estimate the salvage value of Method 2 will be 26% higher after three years than it is after five years. If the MARR is 12% per year, which method should the company select?

The company should select
Answer #1

MARR = 12%

Study period = 3 years

Method 1

Initial cost = 520000

AOC = 230000

Salvage value = 155000

Life = 3 years

Present worth of method 1 = -520000 - 230000* (P/A, 12%, 3) + 155000 *(P/F,12%,3)

= -520000 - 230000*2.4018312 + 155000 *0.71178024

= -962095.25

Method 2

Initial cost = 950000

AOC = 155000

Salvage value = 210000

Life = 5 years

As study period is three year and it is given salvage value will be 26% higher in EOY 3 than at EOY5

So, salvage value at EOY 3 = 1.26 * 210000 = 264600

Present worth of method 2 = -950000 - 155000* (P/A, 12%, 3) + 264600 *(P/F,12%,3)

= -950000 - 155000*2.4018312 + 264600 *0.71178024

= - 1133946.79

Method 1 should be selected as it has less present cost